what is the sum of the solution |3-6x|=15
Answer by mathbird(71)

|3-6x|=15
3-6x=15:3-6x=-15
-6x=15-3:-6x=-15-3
-6x=12:-6x=-18
x=-2:x=3
solution is-2,3
sum of solution=-2+3=1